OpenStudy (anonymous):

find (g o f)(x) for the pair of functions and five the domain. F(x)=7/x, g(x)=3-x

OpenStudy (anonymous):

Use f(x) as input into the function g(x).

OpenStudy (anonymous):

meant to type give the domain

OpenStudy (anonymous):

so g(x) = 3 - 7/x?

Parth (parthkohli):

\[(g \circ f)(x) \implies g(f(x))\]Here,\[(g \circ f)(x) \implies g(f(x)) \implies g\left({7 \over x}\right)\implies 3 - {\left(7 \over x \right)}\]

Parth (parthkohli):

we might actually simplify it as the following:\[{3x \over x} - {7 \over x}\]\[\implies {3x -7 \over x}\]

OpenStudy (anonymous):

Now to find the domain, what is X not allowed to be?

OpenStudy (anonymous):

I have no idea......

OpenStudy (anonymous):

What would break an important math rule and make things go all hay-wire?

OpenStudy (anonymous):

dividing by 0?

OpenStudy (anonymous):

Look where X is. What might happen if it were some particular number?

OpenStudy (anonymous):

Exactly! Don't let X do that!

OpenStudy (anonymous):

so my domain is 0?

OpenStudy (anonymous):

No it is x≠0.

OpenStudy (anonymous):

Or more specifically, it is the set of all numbers except x=0.

OpenStudy (anonymous):

ok. thank you

OpenStudy (anonymous):

so domain is x\[x \neq 0 \]

OpenStudy (anonymous):

In compact math symbols, it is x∈{R:x≠0}, but you can usually just put a " ,x≠0" after your function to state that.

OpenStudy (anonymous):

ok. thank you

Parth (parthkohli):

lol, hay-wire :)

OpenStudy (anonymous):

;-)

OpenStudy (anonymous):

ok so i got another one that says f(x)=x-1where i find (f o f)(x) I came out with (f o f)(x)=x-2 is that right? and how would i find the domain?